OpenStudy (risingtotop):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. (3 points) Equation #1 Equation #2 |x - 4| = 6 |3x + 12| = 18 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to equations show

OpenStudy (risingtotop):

i think its A

OpenStudy (risingtotop):

@Austin.L

OpenStudy (risingtotop):

@quickstudent

OpenStudy (risingtotop):

am i correct?

OpenStudy (austin.l):

With these, you're always going to have 2 solutions to my knowledge.

OpenStudy (risingtotop):

yes i already did them

OpenStudy (austin.l):

Then, if both equations have 2 solutions, that would be correct.. yes?

OpenStudy (risingtotop):

well you have to the the math lol and then tell me if its right -.-

OpenStudy (risingtotop):

you went to engineering school how do you not know how to do this?

OpenStudy (herpderp):

Rude much?

OpenStudy (herpderp):

He's trying to help you, and literally you're being mean ;___; he has a hangover too

OpenStudy (austin.l):

Okay, you said you solved them and got the answers. And you said that you think it is the correct answer. I can work them out for you though. \[|x - 4| = 6\] so \[x - 4 = 6\] & \[x - 4 = -6\] So for this, \[x = 10 \text{ and } x = -2\] \[|3x + 12| = 18\] \[3x + 12 = 18\] \[3x+ 12 = -18\] therefore \[x = -10 \text{ and } x=2\] Both have 2 answers, so therefore, the first answer choice would be correct.

OpenStudy (risingtotop):

okay i got the same thing sorry for the wait